OpenStudy (anonymous):

Solve for the variable. -0.5(r - 2) = -r - 2

OpenStudy (anonymous):

\[-0.5r+1=-r-2 \] \[-0.5r+r =-2-1\] \[.5r=-3\] \[r =3/.5\] \[r =.6\]

OpenStudy (anonymous):

-0.5(r - 2) = -r - 2 -0.5r + 1 = -r - 2 -0.5r + 1 + r = - 2 +0.5r + 1 = - 2 +0.5r = - 2 - 1 +0.5r = - 3 r = - 3/ 0.5 r = - 6